Solve the equation:
4(n-2)-6(n+3)=-14

Answer by stanbon(75887)

4(n-2)-6(n+3)=-14
------
4n - 8 -6n - 18 = -14
----
-2n - 26 = -14
-2n = 12
n = -6
=========================
Cheers,
Stan H.